A vector A ⃗ makes an angle of 20° and B ⃗ makes an angle of 110° with the X-axis. The magnitudes of these vectors are 3 m and 4 m respectively. Find the resultant.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. As shown in the figure, The angle between A ⃗ and B ⃗ = 110° - 20° = 90° |A ⃗ | = 3 and |B ⃗ | = 4m Resultant R = √(A^2+B^2+2AB cos θ) = 5 m Let β be the angle between R ⃗ and A ⃗ Β = tan-1 ((4 sin⁡〖90°〗)/(3+4 cos 90°)) = tan-1 (4/3) = 53° ∴ Resultant vector makes angle (53° + 20°) = 73° with x-axis.